OpenStudy (anonymous):

5-h=2gk+4 solve for g

Parth (parthkohli):

You wanna get \(\rm g\) by itself, right? So first you subtract 4 from both sides. Then you divide both sides by 2k.

OpenStudy (anonymous):

1-h = 2gk (1-h)/2 = gk g= (1-h)/2 • g